State of Econy Probability that Scale will crear n Retum in slock 1 Return on Ru stock a Recession 0.4 10% Boom 8.8% ExpectedUse the following information to complete the question; compute the expected return, the variance and standard deviation for the portfolio.
Answer #1

Expected return in Recession =0.20*10%+0.50*15%+0.30*20% =15.50%
Expected return in Boom =0.20*8%+0.50*4%+0.30*0% =3.60%
Expected Return of Portfolio =0.4*15.50%+0.6*3.60% =8.36%
Variance =0.4*(15.50%-8.36%)^2+0.6*(3.60%-8.36%)^2 =0.3399%
Standard Deviation =Variance^0.5 =0.3399%^0.5=5.83%